a gaseous hydrocarbon gives upon combustion .72gm of water and 3.08gm of carbon dioxide.then what is the empirical formula of hydrocarbon is
1
Expert's answer
2013-06-10T11:07:48-0400
Hydrocarbons react with oxygen according to followingequation:

CxHy + zO2 = xCO2 + y/2H2O
We can find the empirical formula of the hydrocarbon fromthe next ratio:
(x/y) = (m(CO2)/M(CO2)) : (m(H2O)/(M(H2O)/2)) = (3.08 g / 44 g/mol ) : (0.72 g / (18 g/mol / 2 )) = 0.07 : 0.08 = 7 : 8
Thus, the empirical formula of the hydrocarbon isC7H8.